# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

### v2

@@ A.2 Performance of ULTRA
A.3 Performance of Other Potential Shortcuts
In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut. A natural question is whether this is true for just PPR, or if other shortcuts exist. Another bias discussed in KG literature is degree bias [38]. In their study, they show that that KG methods tend to perform better on entities with a higher degree. Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”). Based on this, we study whether the tail degree can is a good predictor of performance.
Table 11: Performance of Tail Degree [38], PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
@@ Appendix B Datasets
B.1 Existing Datasets
We detail the statistics of all existing transductive and inductive datasets in Tables 12, 13, and 14, respectively. We further include the licenses for each in Table 15. Note that we omit YAGO3-10 [39] as [40] show that the dataset is dominated by two duplicate relations, making most triples trivial to classify. Also, we omit NELL-995 and any inductive datasets derived from it due to findings by [31] that show that most triples in the dataset are either too generic or meaningless.
Table 12: Statistics for Transductive Datasets.
@@ B.1 Existing Datasets
FB15k-237 [33] 14,541 237 272,115 17,535 20,466
WN18RR [27] 40,943 11 86,835 3,034 3,134
CoDEx-S [31] 2,034 42 32,888 1,827 1,828
CoDEx-M [31] 17,050 51 185,584 10,310 10,311
CoDEx-L [31] 77,951 69 551,193 30,622 30,622
HetioNet [32] 45,158 24 2,025,177 112,510 112,510
DBpedia100k [37] 99,604 470 597,572 50,000 50,000
Table 13: Statistics for (E) Inductive Datasets.
@@
FB15k-237 v1 [7] 180 1,594 4,245 1,594 4,245 489 1,093 1,993 205
FB15k-237 v2 [7] 200 2,608 9,739 2,608 9,739 1,166 1,660 4,145 478
FB15k-237 v3 [7] 215 3,668 17,986 3,668 17,986 2,194 2,501 7,406 865
FB15k-237 v4 [7] 219 4,707 27,203 4,707 27,203 3,352 3,051 11,714 1,424
WN18RR v1 [7] 9 2,746 5,410 2,746 5,410 630 922 1,618 188
WN18RR v2 [7] 10 6,954 15,262 6,954 15,262 1,838 2,757 4,011 441
WN18RR v3 [7] 11 12,078 25,901 12,078 25,901 3,097 5,084 6,327 605
WN18RR v4 [7] 9 3,861 7,940 3,861 7,940 934 7,084 12,334 1,429
ILPC-S [10] 48 10,230 78,616 6,653 20,960 2,908 6,653 20,960 2902
ILPC-L [10] 65 46,626 202,446 29,246 77,044 10,179 29,246 77,044 10,184
Table 14: Statistics for (E, R) Inductive Datasets.
@@
FB-25 [9] 5,190 163 91,571 4,097 216 17,147 5,716 4,097 216 17,147 5,716
FB-50 [9] 5,190 153 85,375 4,445 205 11,636 3,879 4,445 205 11,636 3,879
FB-75 [9] 4,659 134 62,809 2,792 186 9,316 3,106 2,792 186 9,316 3,106
FB-100 [9] 4,659 134 62,809 2,624 77 6,987 2,329 2,624 77 6,987 2,329
WK-25 [9] 12,659 47 41,873 3,228 74 3,391 1,130 3,228 74 3,391 1,131
WK-50 [9] 12,022 72 82,481 9,328 93 9,672 3,224 9,328 93 9,672 3,225
WK-75 [9] 6,853 52 28,741 2,722 65 3,430 1,143 2,722 65 3,430 1,144
WK-100 [9] 9,784 67 49,875 12,136 37 13,487 4,496 12,136 37 13,487 4,496
Table 15: Dataset Licenses.
Datasets License
FB15k-237 CC-BY-4.0
WN18RR MIT License
CoDEx-S/M/L MIT License
HetioNet CC0 1.0 Universal
DBpedia100k None
FB15k-237 v1/v2/v3/v4 None
WN18RR v1/v2/v3/v4 None
ILPC-S/L MIT License
FB-25/50/75/100 None
WK-25/50/75/100 None
B.2 New Datasets
We further detail the statistics of all the new datasets in Tables 16 and 17.
Table 16: Statistics for New (E) Inductive Datasets.
@@ B.2 New Datasets
Inference 2 37,927 2455 12 4,187 0.19
Table 17: Statistics for New (E, R) Inductive Datasets.
@@
Inference 2 8,821 2,606 28 979 56.8% 0.60
Appendix C Inductive Generation Experiments
We give more details on the experiments conducted in Section 3.3 and shown in Figures 4. Given the transductive dataset FB15k-237 [33], we generate a number of different inductive datasets using the common procedure used by [7, 9]. We describe this procedure in detail in Section 2. Note that for simplicity, we only created datasets for the (E) inductive task. However, the same conclusions should hold for (E, R). Lastly, we only create one graph for inference as is done in previous work.
The default neighborhood size is set to 50 for the inductive datasets created by [7] and [9]. For [9] they consider 10 and 20, initial entities for train and inference, respectively. This information is not reported for [7].
To simulate the effect of the different parameter values, we create inductive datasets by using a number of different combinations of parameters. For each parameter, we modify it while holding the others constant. This allows us to explore the effect of just that variable. It allows us to avoid running an excessive amount of experiments. The default values are those used in [9] and noted above. In total, there are 17 parameter configurations. For each parameter configuration, we create 3 different datasets through the use of 3 different random seeds. This results in a total of 51 inductive datasets. We then evaluate the performance of PPR on each inference graph and calculate the $\Delta\text{SPD}$. The results for those datasets with the same configuration are then averaged together. All possible configurations and their resulting mean statistics can be found in Table 18. These results are visualized in the main content in Figures 4.
Table 18: FB15k-237 - Inductive Generative Experiments Configurations. Results are over 3 Random Seeds.
@@ Appendix C Inductive Generation Experiments
Appendix D Additional Experimental Settings
Training Settings: Each model was trained on either a: Tesla V100 32Gb, NVIDIA RTX A6000 48Gb, NVIDIA RTX A5000 24Gb, or Quadro RTX 8000 48Gb. All models were implemented in PyTorch [41] and Torch-Geometric [42].
Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method. For each method, we tried to follow the recommended ranges used by the authors. For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$. For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$. For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$. In their experiments, NBFNet [8] uses the same hyperparameters across all datasets. We therefore do the same. For Neural LP [13] we also use the default set of hyperparameters that is shared across datasets.
Evaluation: For a given test triple $(s,r,o)$, we strive to predict both $s$ and $o$ individually. This is framed as a ranking problem where we want the probability of the correct entity (e.g., $o$) to rank higher than all possible negative entities. We can calculate the rank of the true entity over all negatives by the following where $p(\cdot)$ denotes the model probability for a triple:
[…]
We note that a lower rank, indicates better performance. In practice, some of the negative entities will be true triples observed during training. As such, following [4] we use the filtered setting and remove those entities from the ranking. Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute the Hits@K%. Hits@Kcomputes the percentage of sample where the following inequality holds true, $\text{Rank}\leq K$.
Appendix E ULTRA Settings
When evaluating ULTRA [34] we use the 0-shot setting. By default, we use the checkpoint they provided that was trained on the transductive datasets: FB15k-237 [33], WN18RR [27], CoDEx-M [31], and NELL-995 [43]. However, since we create our own inductive splits from FB15k-237, CoDEx-M, and WN18RR, there is risk of test leakage. That is, some triples that may have been in the transductive training graph, may now be a test sample in one of our splits. This gives ULTRA an unfair advantage as it has already seen that triple.
@@ Appendix E ULTRA Settings
The results are shown in Table 19, where the “W/o” column represents the results when removing the parent transductive dataset and “With” corresponds to the original pre-trained checkpoint provided by the authors of ULTRA [34]. In practice, we observe a small but noteworthy decrease in performance when removing the offending transductive dataset from the pre-trained model.
Table 19: Performance (Hits@10) of ULTRA when pre-trained and w/o the parent transductive dataset. Note that this doesn’t apply to HetioNet.
@@
Appendix F Implementation
All code used for experiments in this study can be found in the following anonymized repository – https://github.com/HarryShomer/Better-Inductive-KGC. Please see the README in the repository for more information on how to run the different experiments.
Appendix G Limitations
One potential limitation of our study is that the new inductive datasets are still being created from existing transductive datasets. While our method can help create better and more realistic inductive scenarios, this is limiting as we are still reliant on good quality transductive datasets. Future work can focus on creating inductive KGC datasets directly from existing large-scale knowledge graphs, thereby bypassing the need to sample from existing transductive datasets.
Appendix H Impact Statement
Our method contributed positively to the field of knowledge graph reasoning. By introducing newer and better inductive datasets, we better align the task of KGC to it’s real-world applications. This is essential, because if we want to perform inductive KGC in real-world tasks, we have to be certain that the methods actually work well. As such, the new datasets can serve as a stronger barometer of actual performance on inductive KGC and will help spur the development of newer and more effective KGC techniques. Since KGC has applications in many different fields including question answering, biology, and recommender systems, there is a strong need for methods that can perform well. We have also carefully considered the broader impact of our work from different perspectives and find that there is no apparent risk.

### v3

@@ A.2. Performance of ULTRA
Table 11. Performance of Tail Degree (Shomer et al., 2023a), PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
@@
A.3. MRR Performance
We further include the performance on the new datasets when evaluating via MRR. The results are shown in Tables 12 and 13 for the (E) and (E, R) results, respectively. We can observe that these results are highly consistent when evaluating via Hits@10. That is, under the (E) setting both NBFNet and RED-GNN perform best, while under the (E, R) setting DEq-InGram also excels. This suggests that our observations hold under multiple evaluation metrics.
Table 12. (E) Inductive Results (MRR) for supervised methods.
Models CoDEx-M WN18RR HetioNet
Inference 1 Inference 1 Inference 2 Inference 1 Inference 2
PPR 3.3 25.6 7.9 2.3 2.1
Neural LP 9.2 ± 11.2 23.9 ± 2.0 10.5 ± 1.8 7.3 ± 9.9 6.7 ± 9.3
NodePiece 3.1 ± 0.3 20.6 ± 0.8 3.3 ± 0.4 5.1 ± 0.6 7.2 ± 0.6
InGram 10.1 ± 1.5 15.9 ± 0.7 3.8 ± 1.1 11.2 ± 0.5 11.5 ± 2.1
DEq-InGram 11.2 ± 0.6 25.4 ± 0.5 9.8 ± 1.7 16.5 ± 2.9 16.9 ± 2.0
RED-GNN 22.1 ± 1.8 67.8 ± 0.5 16.4 ± 1.3 53.8 ± 2.5 68.6 ± 4.3
NBFNet 27.6 ± 0.2 66.6 ± 0.1 17.6 ± 0.1 39.1 ± 1.2 45.2 ± 0.5
Table 13. (E, R) Inductive Results (MRR) for supervised methods. The % of new relations are in parentheses.
Models FB15k-237 CoDEx-M
PPR 3.8 4.7 3.3 7.1
Neural LP 8.8 ± 4.9 11.3 ± 6.6 13.8 ± 18.8 7.4 ± 9.4
NodePiece 1.4 ± 0.1 2.6 ± 0.2 1.5 ± 0.2 1.6 ± 0.4
InGram 11.2 ± 1.9 9.2 ± 0.9 10.3 ± 1.2 7.2 ± 4.2
DEq-InGram 19.5 ± 2.0 13.4 ± 1.2 18.7 ± 7.6 13.0 ± 1.3
RED-GNN 11.8 ± 2.8 19.8 ± 2.2 16.8 ± 1.6 14.7 ± 5.3
NBFNet 15.8 ± 0.9 15.9 ± 0.2 27.9 ± 8.8 8.5 ± 4.8
A.4. Performance of Other Potential Shortcuts
In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut. A natural question is whether this is true for just PPR, or if other shortcuts exist. Another bias discussed in KG literature is degree bias (Shomer et al., 2023a). In their study, they show that that KG methods tend to perform better on entities with a higher degree. Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”). Based on this, we study whether the tail degree can is a good predictor of performance.
@@ Appendix B Datasets
B.1. Existing Datasets
We detail the statistics of all existing transductive and inductive datasets in Tables 14, 15, and 16, respectively. Note that we omit YAGO3-10 (Mahdisoltani et al., 2013) as (Akrami et al., 2020) show that the dataset is dominated by two duplicate relations, making most triples trivial to classify. Also, we omit NELL-995 and any inductive datasets derived from it due to findings by (Safavi and Koutra, 2020) that show that most triples in the dataset are either too generic or meaningless.
Table 14. Statistics for Transductive Datasets.
@@ B.1. Existing Datasets
FB15k-237 (Toutanova and Chen, 2015) 14,541 237 272,115 17,535 20,466
WN18RR (Dettmers et al., 2018) 40,943 11 86,835 3,034 3,134
CoDEx-S (Safavi and Koutra, 2020) 2,034 42 32,888 1,827 1,828
CoDEx-M (Safavi and Koutra, 2020) 17,050 51 185,584 10,310 10,311
CoDEx-L (Safavi and Koutra, 2020) 77,951 69 551,193 30,622 30,622
HetioNet (Himmelstein et al., 2017) 45,158 24 2,025,177 112,510 112,510
DBpedia100k (Ding et al., 2018) 99,604 470 597,572 50,000 50,000
Table 15. Statistics for (E) Inductive Datasets.
@@
FB15k-237 v1 (Teru et al., 2020) 180 1,594 4,245 1,594 4,245 489 1,093 1,993 205
FB15k-237 v2 (Teru et al., 2020) 200 2,608 9,739 2,608 9,739 1,166 1,660 4,145 478
FB15k-237 v3 (Teru et al., 2020) 215 3,668 17,986 3,668 17,986 2,194 2,501 7,406 865
FB15k-237 v4 (Teru et al., 2020) 219 4,707 27,203 4,707 27,203 3,352 3,051 11,714 1,424
WN18RR v1 (Teru et al., 2020) 9 2,746 5,410 2,746 5,410 630 922 1,618 188
WN18RR v2 (Teru et al., 2020) 10 6,954 15,262 6,954 15,262 1,838 2,757 4,011 441
WN18RR v3 (Teru et al., 2020) 11 12,078 25,901 12,078 25,901 3,097 5,084 6,327 605
WN18RR v4 (Teru et al., 2020) 9 3,861 7,940 3,861 7,940 934 7,084 12,334 1,429
ILPC-S (Galkin et al., 2022) 48 10,230 78,616 6,653 20,960 2,908 6,653 20,960 2902
ILPC-L (Galkin et al., 2022) 65 46,626 202,446 29,246 77,044 10,179 29,246 77,044 10,184
Table 16. Statistics for (E, R) Inductive Datasets.
@@
FB-25 (Lee et al., 2023) 5,190 163 91,571 4,097 216 17,147 5,716 4,097 216 17,147 5,716
FB-50 (Lee et al., 2023) 5,190 153 85,375 4,445 205 11,636 3,879 4,445 205 11,636 3,879
FB-75 (Lee et al., 2023) 4,659 134 62,809 2,792 186 9,316 3,106 2,792 186 9,316 3,106
FB-100 (Lee et al., 2023) 4,659 134 62,809 2,624 77 6,987 2,329 2,624 77 6,987 2,329
WK-25 (Lee et al., 2023) 12,659 47 41,873 3,228 74 3,391 1,130 3,228 74 3,391 1,131
WK-50 (Lee et al., 2023) 12,022 72 82,481 9,328 93 9,672 3,224 9,328 93 9,672 3,225
WK-75 (Lee et al., 2023) 6,853 52 28,741 2,722 65 3,430 1,143 2,722 65 3,430 1,144
WK-100 (Lee et al., 2023) 9,784 67 49,875 12,136 37 13,487 4,496 12,136 37 13,487 4,496
B.2. New Datasets
We further detail the statistics of all the new datasets in Tables 17 and 18.
Table 17. Statistics for New (E) Inductive Datasets.
@@ B.2. New Datasets
Inference 2 37,927 2,455 12 4,187 0.19
Table 18. Statistics for New (E, R) Inductive Datasets.
@@
Inference 2 8,821 474 28 979 56.8% 0.60
Table 19. FB15k-237 - Inductive Generative Experiments Configurations. Results are over 3 Random Seeds.
@@
Appendix C Inductive Generation Experiments
We give more details on the experiments conducted in Section 3.3 and shown in Figures 4. Given the transductive dataset FB15k-237 (Toutanova and Chen, 2015), we generate a number of different inductive datasets using the common procedure used by (Teru et al., 2020; Lee et al., 2023). We describe this procedure in detail in Section 2. Note that for simplicity, we only created datasets for the (E) inductive task. However, the same conclusions should hold for (E, R). Lastly, we only create one graph for inference as is done in previous work.
The default neighborhood size is set to 50 for the inductive datasets created by (Teru et al., 2020) and (Lee et al., 2023). For (Lee et al., 2023) they consider 10 and 20, initial entities for train and inference, respectively. This information is not reported for (Teru et al., 2020).
To simulate the effect of the different parameter values, we create inductive datasets by using a number of different combinations of parameters. For each parameter, we modify it while holding the others constant. This allows us to explore the effect of just that variable. It allows us to avoid running an excessive amount of experiments. The default values are those used in (Lee et al., 2023) and noted above. In total, there are 17 parameter configurations. For each parameter configuration, we create 3 different datasets through the use of 3 different random seeds. This results in a total of 51 inductive datasets. We then evaluate the performance of PPR on each inference graph and calculate the $\Delta\text{SPD}$. The results for those datasets with the same configuration are then averaged together. All possible configurations and their resulting mean statistics can be found in Table 19. These results are visualized in the main content in Figures 4.
Appendix D ULTRA Settings
When evaluating ULTRA (Galkin et al., 2023) we use the 0-shot setting. By default, we use the checkpoint they provided that was trained on the transductive datasets: FB15k-237 (Toutanova and Chen, 2015), WN18RR (Dettmers et al., 2018), CoDEx-M (Safavi and Koutra, 2020), and NELL-995 (Xiong et al., 2017). However, since we create our own inductive splits from FB15k-237, CoDEx-M, and WN18RR, there is risk of test leakage. That is, some triples that may have been in the transductive training graph, may now be a test sample in one of our splits. This gives ULTRA an unfair advantage as it has already seen that triple.
@@ Appendix D ULTRA Settings
The results are shown in Table 22, where the “W/o” column represents the results when removing the parent transductive dataset and “With” corresponds to the original pre-trained checkpoint provided by the authors of ULTRA (Galkin et al., 2023). In practice, we observe a small but noteworthy decrease in performance when removing the offending transductive dataset from the pre-trained model.
Table 20. PPR Performance on PediaTypes (Gao et al., 2023) datasets.
Dataset 50 Negatives All Negatives
Neural Hits@10 PPR Hits@10 % Difference PPR Hits@10 $\Delta\text{SPD}$
EN-FR 95.4% 96.0% -0.6% 44.1% 2.7
FR-EN 90.6% 96.2% -5.8% 48.1% 2.3
EN-DE 97.7% 93.9% 4.0% 34.7% 2.1
DE-EN 95.0% 90.6% 4.9% 20.8% 1.6
DB-WD 86.6% 76.4% 13.4% 23.5% 0.6
WD-DB 91.5% 92.0% -0.5% 42.9% 2.5
DB-YG 71.5% 76.9% -7.0% 15.2% 2.7
YG-DB 80.5% 85.1% -5.4% 30.7% 3.5
Table 21. PPR Performance on WikiTopics (Gao et al., 2023) datasets.
Dataset 50 Negatives All Negatives
Neural Hits@10 PPR Hits@10 % Difference PPR Hits@10 $\Delta\text{SPD}$
Art 92.3% 79.3% 16.4% 14.7% 1.1
Award 93.6% 95.8% -2.3% 24.2% 3.4
Edu 97.0% 72.4% 34.0% 20.8% 2.7
Health 98.9% 91.5% 8.1% 59.6% 3.3
Infra 97.8% 97.2% 0.6% 61.4% 5.4
Sci 96.7% 84.3% 14.7% 25.5% 4.4
Sport 81.2% 82.9% -2.1% 14.4% 0.3
Tax 93.7% 64.1% 46.1% 29.0% 0.8
Loc N/A 96.5% N/A 74.1% 1.4
Org N/A 94.0% N/A 8.9% 2.8
People N/A 94.9% N/A 64.6% 2.7
Table 22. Performance (Hits@10) of ULTRA when pre-trained and w/o the parent transductive dataset. Note that this doesn’t apply to HetioNet.
@@
Appendix E Analysis of the WikiTopics and PediaTypes Datasets
We further include an analysis of the WikiTopics and PediaTypes inductive datasets introduced by Gao et al. (2023). For these datasets, the train and inference graphs contain completely separate entities and relations. For each dataset we record the (a) Neural SOTA performance (taken from (Gao et al., 2023), (b) PPR performance, (c) % difference in performance, and (d) $\Delta\text{SPD}$. In their original study, Gao et al. (2023) use 50 random negatives during evaluation, instead of all negative entities used in ours. To account for this, we compute the PPR performance using both evaluation settings. The results for the PediaTypes and WikiTopics datasets are in Tables 20 and 21, respectively.
From the results we can make multiple observations. First, PPR performs well under both settings. Under the all negatives setting, the mean PPR performance is 32.5% and 38.3% on PediaTypes and WikiTopics, respectively. Second, under 50 negatives, there is little difference in the PPR and Neural performance. Specifically, the % difference in performance is 0.4% and 14.2%, respectively, on PediaTypes and WikiTopics. This suggests that neural methods can only slightly outperform PPR under this setting. Lastly, the $\Delta\text{SPD}$ is high on both sets of datasets. This further corroborates our findings in Section 3.
This analysis shows that the WikiTopics and PediaTopics datasets introduced by Gao et al. (2023) fall prey to the same shortcut found in other inductive datasets studied in Section 3.